Sandhill Company purchased equipment for $236,430 on October 1, 2017. It is estimated that the equipment will have a useful life of 8 years and a salvage value of $13,320. Estimated production is 40,200 units and estimated working hours are 20,000. During 2017, Sandhill uses the equipment for 530 hours and the equipment produces 1,000 units.

Compute depreciation expense under each of the following methods. Sandhill is on a calendar-year basis ending December 31. (Round rate per hour and rate per unit to 2 decimal places, e.g. 5.35 and final answers to 0 decimal places, e.g. 45,892.)

(a) Straight-line method for 2017 $

(b) Activity method (units of output) for 2017 $

(c) Activity method (working hours) for 2017 $

(d) Sum-of-the-years'-digits method for 2019 $

(e) Double-declining-balance method for 2018 $
Answer #1

A) straight line method

Depreciation = cost- salvage value / useful life

= 236,430 - 13,320 / 8

= 27,889 per year

Depreciation for 2017 = 27,889 x 3/12 = 6,972

B) Activity method (units of output)

Depreciation rate = cost- salvage value / useful life in units

= 236,430 - 13,320 / 40,200

   = 5.55 per unit

Depreciation for 2017 = 1,000 x 5.55 = 5,550

C) Activity method ( working hour)

Depreciation rate = cost- salvage value / useful life in hours

= 236,430 - 13,320 / 20,000

   = 11.15 per hour

Depreciation for 2017 = 530 x 11.15 = 5,910

D) sum of the year digit method

1+2+3+4+5+6+7+8 = 36

Depreciation for second year = (cost - salvage value) x 7/36

= (236,430 - 13,320) x 7/36

= 43,382

Depreciation for third year = (cost - salvage value) x 6/36

= (236,430 - 13,320) x 6/36

= 37,185

Depreciation from 1, january 2019 to 30, september 2019 = 43,382 x 9/12 = 32,536

Depreciation from 1, october 2019 to 31, december 2019 = 37,185 x 3 /12 = 9,296

Total depreciation of 2019 = 41,832

E) double declining balance method

Depreciation rate = 100 / useful life x 2

= 100 / 8 x 2

= 25%

Depreciation for first year = 236,430 x 25% = 59,107

Depreciation for second year = 177,323*x 25% = 44,330

* 236,430 - 59,107 = 177,323

Depreciation from 1, january 2018 to 30, september 2018 = 59,107 x 9/12 = 44,330

Depreciation from 1, october 2018 to 31, december 2018 = 44,330 x 3 /12 = 11,082

Total depreciation for 2018 = 55,412